Q)The value of the parameter a,for which the function f(x)=1+ax,a is not equal to zero is the inverse of itself,is 
a)-2  b)-1  c)1  d)2
PLZ solve it sir

Grade:11

1 Answers

Vikas TU
7 years ago
Inverse of f(x) is:
y = ax + 1
ax = y – 1
x = (y-1)/a
or
y(x) = (x-1)/a
then,
(x-1)/a = (1+ax)
solve the quadratic and fin the dicriminant and its range.